A route up a mountain is 20 km long john followed this routeat the average speed of x km/h
a)Write down an expression in trems of x for a number of hourshe tooks to walk up the mountain
b)he came down the mountain by a different route the lenght ofthis route was 25 km his average speed coming down the mountain was2 m per hour greater then his average speed going up the mountainwrite down an expression in terms of x for the number of hours hetook to walk down
c)if took john one and half hours less to come downthen to go up write down an equation in x and show that itsimplifies to 3x^2+16x-80-0

Answer by Boreal(15235)

speed*time=distance
x*time=20
time=20/x
==========
(x+2)*time=25
time=25/(x+2)
==========
The second time is 1.5 less than the first
Therefore,(20/x)-1.5=(25/(x+2))
multiply everything by x*(x+2)
20(x+2)-1.5(x^2+2x)=25x
20x+40-1.5x^2-3x=25x
0=1.5x^2-8x-40
3x^2-16x-80=0